radiation force due to source of power p on perfectly reflecting surface?

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
E = m0c2

or , m0c = E/c

Change in momentum Δp = 2m0c = 2E/c

F = dp/dt = 2p/c (As dE/dt =p)